OpenStudy (anonymous):

Solve the system of equations below algebraically 5y + 3x = -10 \[-\frac{ 5 }{ 3 }y = x - \frac{ 4 }{ 3 }\]

jimthompson5910 (jim_thompson5910):

First I'd start off by multiplying both sides of equation two by the LCD 3 to go from \[\Large -\frac{ 5 }{ 3 }y = x - \frac{ 4 }{ 3 }\] to \[\Large -5y = 3x - 3\]

jimthompson5910 (jim_thompson5910):

Then subtract 3x from both sides to get \[\Large -5y = 3x - 3\] \[\Large -5y - 3x = - 3\] \[\Large -3x - 5y = - 3\]

jimthompson5910 (jim_thompson5910):

So you now have the system 5y + 3x = -10 -3x - 5y = -3 which is really the same thing as this system 3x + 5y = -10 -3x - 5y = -3 I'll let you finish up